# SYM3D: Learning Symmetric Triplanes for Better 3D-Awareness of GANs

### v1

@@ Abstract
Despite the growing success of 3D-aware GANs, which can be trained on 2D images to generate high-quality 3D assets, they still rely on multi-view images with camera annotations to synthesize sufficient details from all viewing directions. However, the scarce availability of calibrated multi-view image datasets, especially in comparison to single-view images, has limited the potential of 3D GANs. Moreover, while bypassing camera pose annotations with a camera distribution constraint reduces dependence on exact camera parameters, it still struggles to generate a consistent orientation of 3D assets. To this end, we propose SYM3D, a novel 3D-aware GAN designed to leverage the prevalent reflectional symmetry structure found in natural and man-made objects, alongside a proposed view-aware spatial attention mechanism in learning the 3D representation. We evaluate SYM3D on both synthetic (ShapeNet Chairs, Cars, and Airplanes) and real-world datasets (ABO-Chair), demonstrating its superior performance in capturing detailed geometry and texture, even when trained on only single-view images. Finally, we demonstrate the effectiveness of incorporating symmetry regularization in helping reduce artifacts in the modeling of 3D assets in the text-to-3D task.
@@ 1 Introduction
Despite their impressive performance, current 3D GANs face several challenges. Multi-view images remain indispensable for synthesizing 3D assets with plausible details from all viewing angles [10, 21, 7]. While it is possible to train a 3D GAN using only single-view images, the quality declines substantially. Meanwhile, 3D GANs without camera annotations as supervision often produce 3D assets with arbitrary orientations, posing difficulties for users in accurately positioning these assets in downstream applications, due to the absence of a consistent local coordinate frame. Furthermore, learning objects in a consistent orientation allows the model to exploit structural information across examples. Restricting the sampling of camera locations during the rendering stage of training to align with the camera distribution in the training dataset does not fully resolve this issue [18, 10]. Given the scarcity of multi-view datasets with accurate camera annotations in real-world applications, this necessitates an approach capable of learning from single-image settings, where each object offers only one view for training, with only camera distribution available.
@@ 2 Related work
3D-aware GANs with Triplane Representation Since EG3D [4] introduces the concept of triplane representation, this representation has become increasingly popular in 3D modeling for its efficiency and effectiveness. A line of works focuses on improving triplane representations, for example, He et al. [11] inject an extra axis to 2D triplane by preserving information associated with projection distance. Wu and Zheng [24] develop a multi-scale triplane in a hierarchical fashion that enables learning 3D shapes from a single example, progressively refining from coarse to detailed features. Another line of works distills the triplane representation to describe a single object. Trevithick et al. [22] learn a triplane representation to describe an unposed face from a pretrained EG3D, which facilitates real-time, photo-realistic 3D face rendering. Bhattarai et al. [3] improve GAN inversion by learning offsets to adjust the triplane representation. Our work enhances triplane representation from the perspective of structure awareness. This is achieved by integrating view-wise spatial attention and symmetry regularization into triplane learning. It facilitates a more holistic representation of the object in training 3D-aware generators.
Reflectional Symmetry in Image Synthesis Symmetric objects are widespread in nature, architecture, and art, and reflectional symmetry plays a key role in human visual perception [16]. This concept has been extensively utilized in computer vision research. A significant research direction in this domain focuses on applying reflectional symmetry principles to enhance the rendering of 2D images. Wu et al. [25] investigate the inference of 3D deformable objects from single images by employing symmetric structures to distinguish between depth, albedo, viewpoint, and illumination components. NeRD [26] presents a neural detector designed to identify 3D reflection symmetry in objects, estimating the normal vectors of their mirror planes. Yin et al. [28] have advanced 3D GAN inversion techniques by training with mirrored images, leveraging symmetry to enhance the quality of the results. SymmNeRF [15] represents an innovative strategy that explicitly incorporates symmetry into the training of neural radiance fields, utilizing both pixel-aligned image features and their symmetric analogs as additional training inputs. Our work applies symmetry regularization to the geometry and texture representations of an object, which significantly enhances the triplane’s representation ability, offering a more effective representation for creating realistic 3D assets.
@@ 3.3 Reflectional Symmetry Regularization
For the texture triplane, we implement a methodology where each pixel’s features are combined with those of its symmetrical counterpart:
@@
3.4 Training Objectives
$L_{D}$ $=-\mathbb{E}[\log(1-\mathrm{D}(I_{f}))]-\mathbb{E}[\log(D(I_{r}))]+\lambda\mathbb{E}\left[\|\nabla_{I_{r}}D(I_{r})\|^{2}\right],$ (8)
$L_{G}$ $=-\mathbb{E}[\log(D(I_{f}))]+\alpha\mathcal{R}(G)+\beta\mathcal{R}(A),$ (9)
where $I_{r}$ denotes real data while $I_{f}$ denotes data produced by the generator (i.e. RGB image, silhouettes). The third element in Eq 8 represents the gradient penalty, with $\lambda$ indicating the associated weighting coefficient for this term. The terms $\alpha$ and $\beta$ are hyperparameters that control the symmetry regularization on the geometry triplane and the attention maps, respectively. We also incorporate the shape regularization utilized in GET3D to remove internal floating faces. Our discriminator does not require the camera pose of real images, which is inaccessible in most real-world datasets.
4 Experiment
4.1 Settings
Datasets We conduct experiments on the synthetic ShapeNet [6] dataset and the real-world Amazon Berkeley Objects (ABO) [9] dataset. Following GET3D [10], we focus on car and chair categories from ShapeNet, where each object is represented by 24 different views. As suggested by [10], we split each category into training (70%), validation (10 %), and testing (20%) sets. To simulate a more realistic training scenario similar to natural images, our approach only uses a single view of each object in the training set. We define three scenarios based on the range of azimuth angles for the selected view: Scenario 1 (S1) spans 0-360 degrees, Scenario 2 (S2) covers 0-180 degrees, and Scenario 3 (S3) encompasses 0-120 degrees. For S2 and S3, we adopt random view flips to enhance view diversity. For ABO datasets, we run experiments on its chair category, which has 1158 objects. All experiments are conducted at a resolution of $1024\times 1024$.
Competitors To the best of our understanding, our method marks the initial attempt to integrate a symmetry prior into a triplane representation. While, GET3D is unique in its approach to separate geometry and texture during the training of a 3D generative model, making it the benchmark in this field. As such, GET3D serves as our primary reference point. Furthermore, we also draw comparisons with another SOTA work in 3D-aware image synthesis, OrthoPlanes for 3D (OP3D) [11], which enhances the triplane representation by maintaining information related to the projection distance.
Implementations To facilitate fair comparative evaluation, we adopt the same setup of [10] including the training configuration as given in its open source code [^1] https://github.com/nv-tlabs/GET3D. Experiments are done on 8 A100 GPUS. In real-world settings, accurately determining camera poses can be difficult. As a result, we adopt a strategy of training the discriminator without camera pose annotation, opting instead for a fixed camera distribution approach, as demonstrated to be effective in previous studies [18, 10]. For all our experiments, we utilize the camera distribution defined in GET3D. It’s important to mention that the camera distribution is not completely covered in S3. For hyper-parameters, we set $\alpha=100$, $\beta=10$ in the experiments.
Metrics To assess the quality of the generated objects, we examine both their geometry and texture. For geometric, we use Coverage (COV) and Minimum Matching Distance (MMD) [1] metrics, which assess the average quality and diversity of the shapes. The comparison between two shapes is using the Chamfer Distance (CD) and the Light Field Distance (LFD) [8], measuring their similarity. For OP3D [11], we use marching cubes to extract the underlying geometry. For texture, we utilize the Fréchet Inception Distance (FID) [12], which is applied to 2D rendered images of the objects.
@@ 4.1 Settings
4.2 Main Results
Quantitative Results The quantitative results for the ShapeNet chairs and cars are presented in Table 1. Our observations are as follows: (1) In comparison to OP3D [11], which utilizes a single triplane representation to encode both shape and texture, GET3D [10] and our SYM3D, which separate shape and texture learning, exhibit improved generation in quality and diversity. (2) Models trained on various splits show that S1 outperforms S2, which in turn outperforms S3. Specifically, S1 encompasses a full 360-degree view of an object category, S2 achieves a 360-degree view through image flipping, while S3 is limited to a 240-degree view. The results further suggest that incorporating views from various angles, including those from other objects, enhances the learning of a 3D representation. (3) All models are trained without conditioning discriminator on camera poses. Our approach, which incorporates symmetry assumed in a canonical space, yields superior results in both COV and MMD metrics. This indicates that the shapes learned by SYM3D closely resemble those in a canonicalized form. (4) SYM3D outperforms GET3D in majority metrics, demonstrating the benefits of integrating symmetry regularization and view-wise spatial attention into 3D representations. Consequently, SYM3D proves to be an effective approach that enhances representation in 3D-aware generation.
@@ 4.2 Main Results
Qualitative Comparisons Figure 4 provides qualitative comparisons against competitors in terms of generated 2D images. In general, SYM3D achieves a more realistic appearance across different settings. In the more challenging settings S2, S3, and real-world ABO-chair, OP3D struggles to generate a complete scene, and GET3D tends to generate distortions.
Figure 5 provides comparisons across various camera positions. GET3D and SYM3D are able to generate view-consistent images across different camera views but OP3D fails to do that. This demonstrates that decoupling shape and texture benefits the 3D-aware generation.
@@ 4.3 Properties of Learned Triplane
View-wise Triplanes We use a similarity matrix to measure difference among feature maps in the geometry triplane. The geometry triplane is a $N\times N\times 3C$ feature tensor, which we flatten to a 2D tensor $3C\times N^{2}$.
@@
Robustness to Biased Views Figure 1 shows the results from models trained on the incomplete view dataset: ShapNet Chair-S3. Overall, SYM3D shows its superiority in generating complete objects.
@@
VSA $\mathcal{R}(G)$ $\mathcal{R}(A)$ $\mathcal{R}(T)$ FID
@@
Table 2: Evaluation of individual components in SYM3D. Experiments are done on ShapeNet Chair-S1 dataset.
Consistent Camera Orientations The fixed camera distribution strategy ensures that all objects produced by both GET3D and SYM3D are uniformly positioned and oriented. By applying symmetry regularization, our method secures a consistent orientation for objects in the chair category with respect to the symmetry plane (Figure 8).
4.4 Further Analysis
The effect of different components We validate the design of our framework by ablating four components using the ShapeNet Chair-S1 dataset. SVE introduces view-wise spatial enhancement in Section 3.2, $\mathcal{R}(G)$ adds symmetry in feature maps as in Eq. 5, $\mathcal{R}(A)$ adds symmetry in attention maps as in Eq. 6, $\mathcal{R}(T)$ supplements a point’s feature with its symmetric counterpart as in Eq. 7. Our findings indicate that each component independently improves the quality of generation, with the collective implementation of all elements resulting in the most effective model.
@@ 4.4 Further Analysis
With window self-attention To show the effectiveness of the proposed attention module, we compare SYM3D against a variant that uses window self-attention [17] across three views, without employing attention symmetry. Considering memory limitations, we chose two window sizes: Wins=4 and Wins=8. Table 3 indicates that SYM3D, while adding negligible increases in parameters and computational cost, results in significant improvements.
On a more complex category For evaluating the generality of proposed method on more complex categories, we test on airplane dataset in ShapeNet [6]. We render 24 views for each object and select one view as training set. As shown in Table 4, we obtain consistent results as on ShapeNet-Chair and ShapeNet-Car in Table 1.
Applied on the text-to-3D task MTN [27] introduces a multi-scale triplane representation for the text-to-3D task. We apply symmetric regularization (see Eq 5) to these multi-scale triplanes. We develop two versions: W/O SYM and W SYM, whose results are in Figure 9. Our observations show that W/O SYM often generates salient artifacts, such as extra or malformed legs and ears when modeling a cat, and a dog with three forelegs. Conversely, W SYM produces cats and dogs without these defects, proving that symmetric priors effectively eliminate such artifacts in 3D model creation.
Figure 9: Comparisons on text-to-3D. W/O SYM produces artifacts in RGB and normal images, while W SYM with symmetry regularization loss avoids this.
@@ 5 Conclusion
Discussion We present SYM3D for learning symmetric triplanes for improving 3D-awareness of GANs when trained on single images without camera pose annotation. While effective it has several limitations that future work should investigate. Global reflectional symmetry is not always satisfied and some objects satisfy other geometric transformations symmetry such as reflections, translations, rotations, or combinations in local parts (i.e. car tires, chair legs). In future work, we plan to extend our approach to create a large real dataset of common object categories and combat the canonicalization issue as in [2] and symmetry issue as in [19].
Broader Impact We advise against using SYM3D in situations where privacy concerns or incorrect recognition might lead to misuse or other harmful consequences. We urge practitioners to meticulously review and remove biases from datasets before training our model, ensuring it fairly represents a diverse range of skin tones, races, and gender identities.

### v2

@@ Abstract
Despite the growing success of 3D-aware GANs, which can be trained on 2D images to generate high-quality 3D assets, they still rely on multi-view images with camera annotations to synthesize sufficient details from all viewing directions. However, the scarce availability of calibrated multi-view image datasets, especially in comparison to single-view images, has limited the potential of 3D GANs. Moreover, while bypassing camera pose annotations with a camera distribution constraint reduces dependence on exact camera parameters, it still struggles to generate a consistent orientation of 3D assets. To this end, we propose SYM3D, a novel 3D-aware GAN designed to leverage the prevalent reflectional symmetry structure found in natural and man-made objects, alongside a proposed view-aware spatial attention mechanism in learning the 3D representation. We evaluate SYM3D on both synthetic (ShapeNet Chairs, Cars, and Airplanes) and real-world datasets (ABO-Chair), demonstrating its superior performance in capturing detailed geometry and texture, even when trained on only single-view images. Finally, we demonstrate the effectiveness of incorporating symmetry regularization in helping reduce artifacts in the modeling of 3D assets in the text-to-3D task. Project is at https://jingyang2017.github.io/sym3d.github.io/
@@ 1 Introduction
Despite their impressive performance, current 3D GANs face several challenges. Multi-view images remain indispensable for synthesizing 3D assets with plausible details from all viewing angles [11, 22, 8]. While it is possible to train a 3D GAN using only single-view images, the quality declines substantially. Meanwhile, 3D GANs without camera annotations as supervision often produce 3D assets with arbitrary orientations, posing difficulties for users in accurately positioning these assets in downstream applications, due to the absence of a consistent local coordinate frame. Furthermore, learning objects in a consistent orientation allows the model to exploit structural information across examples. Restricting the sampling of camera locations during the rendering stage of training to align with the camera distribution in the training dataset does not fully resolve this issue [19, 11]. Given the scarcity of multi-view datasets with accurate camera annotations in real-world applications, this necessitates an approach capable of learning from single-image settings, where each object offers only one view for training, with only camera distribution.
@@ 2 Related work
3D-aware GANs with Triplane Representation. Since EG3D [6] introduces the concept of triplane representation, this representation has become increasingly popular in 3D modeling for its efficiency and effectiveness. A line of works focuses on improving triplane representations, for example, He et al. [12] inject an extra axis to 2D triplane by preserving information associated with projection distance. Wu and Zheng [25] develop a multi-scale triplane in a hierarchical fashion that enables learning 3D shapes from a single example, progressively refining from coarse to detailed features. Another line of works distills the triplane representation to describe a single object. Trevithick et al. [23] learn a triplane representation to describe an unposed face from a pretrained EG3D, which facilitates real-time, photo-realistic 3D face rendering. Bhattarai et al. [4] improve GAN inversion by learning offsets to adjust the triplane representation. Our work improves triplane representation by introducing structure awareness, which enhances the triplane’s ability to capture distinct viewpoints of an object. This enhancement is achieved in conjunction with symmetry regularization, allowing for a more holistic representation in training 3D-aware generators.
Reflectional Symmetry in Image Synthesis. Symmetric objects are widespread in nature, architecture, and art, and reflectional symmetry plays a key role in human visual perception [17]. This concept has been extensively utilized in computer vision research. A significant research direction in this domain focuses on applying reflectional symmetry principles to enhance the rendering of 2D images. Wu et al. [26] investigate the inference of 3D deformable objects from single images by employing symmetric structures to distinguish between depth, albedo, viewpoint, and illumination components. NeRD [27] presents a neural detector designed to identify 3D reflection symmetry in objects, estimating the normal vectors of their mirror planes. Yin et al. [29] have advanced 3D GAN inversion techniques by training with mirrored images, leveraging symmetry to enhance the quality of the results. SymmNeRF [16] represents an innovative strategy that explicitly incorporates symmetry into the training of neural radiance fields, utilizing both pixel-aligned image features and their symmetric analogs as additional training inputs. We introduce symmetry regularization to both geometry and texture, ensuring more accurate and realistic 3D asset generation. Our insight is that the symmetry prior encodes information about the unseen view of an object when only a single view is accessible.
@@ 3.3 Reflectional Symmetry Regularization
While geometric symmetry is more naturally applicable to some categories, texture symmetry can also be beneficial [16]. We implement a methodology where each pixel’s features are combined with those of its symmetrical counterpart:
@@
3.4 Training Objectives
$L_{D}$ $=-\mathbb{E}[\log(1-\mathrm{D}(I_{f}))]-\mathbb{E}[\log(D(I_{r}))]$ (8)
$+\lambda\mathbb{E}\left[\|\nabla_{I_{r}}D(I_{r})\|^{2}\right],$
$L_{G}$ $=-\mathbb{E}[\log(D(I_{f}))]+\alpha\mathcal{R}(G)+\beta\mathcal{R}(A),$
where $I_{r}$ represents real data while $I_{f}$ denotes data produced by the generator (i.e. RGB image, silhouettes). The third element in Eq 8 is the gradient penalty, with $\lambda$ as the weighting coefficient for this term. The hyperparameters $\alpha$ and $\beta$ control the symmetry constraint on the geometry triplane and the attention maps, respectively. We have also adopted the shape regularization utilized in GET3D, which helps eliminate internal floating faces. Our discriminator does not require the camera pose of real images, which is inaccessible in most real-world datasets. Overall, by combining symmetry prior with other losses, our method produces realistic images, not just strictly symmetrical ones, making it suitable for a wide range of categories.
4 Experiment
4.1 Settings
Datasets. We conduct experiments on the synthetic ShapeNet [7] dataset and the real-world Amazon Berkeley Objects (ABO) [10] dataset. Following GET3D [11], we focus on car and chair categories from ShapeNet, where each object is represented by 24 different views. As suggested by [11], we split each category into training (70%), validation (10 %), and testing (20%) sets. To simulate a more realistic training scenario similar to natural images, our approach only uses a single view of each object in the training set. We define three scenarios based on the range of azimuth angles for the selected view: Scenario 1 (S1) spans 0-360 degrees, Scenario 2 (S2) covers 0-180 degrees, and Scenario 3 (S3) encompasses 0-120 degrees. For S2 and S3, we adopt random view flips to enhance view diversity. For ABO datasets, we run experiments on its chair category, which has 1158 objects. All experiments are conducted at a resolution of $1024\times 1024$.
Competitors. To the best of our understanding, SYM3D marks the initial attempt to integrate a symmetry prior into a triplane representation. While, GET3D is unique in its approach to separate geometry and texture during the training of a 3D generative model, making it the benchmark in this field. As such, GET3D serves as our primary reference point. We also draw comparisons with another SOTA work in 3D-aware image synthesis, OrthoPlanes for 3D (OP3D) [12], which enhances the triplane representation by maintaining information related to the projection distance.
@@ 4.1 Settings
Implementations. In real-world settings, accurately determining camera poses can be difficult. As a result, we adopt a strategy of training the discriminator without camera pose condition, opting instead for a fixed camera distribution approach, as demonstrated to be effective in previous studies [19, 11]. For all our experiments, we utilize the camera distribution defined in GET3D. It’s important to mention that the camera distribution is not completely covered in S3. For hyper-parameters, we set $\alpha=100$, $\beta=10$ in the experiments. Additionally, we adopt the same setup of [11] including the training configuration as given in its open source code [^1] https://github.com/nv-tlabs/GET3D. Experiments are done on 8 A100 GPUS.
Metrics. To assess the quality of the generated objects, we examine both their geometry and texture. For geometric, we use Coverage (COV) and Minimum Matching Distance (MMD) [1] metrics, which assess the average quality and diversity of the shapes. The comparison between two shapes is using the Chamfer Distance (CD) and the Light Field Distance (LFD) [9], measuring their similarity. For OP3D [12], we use marching cubes to extract the underlying geometry. For texture, we utilize the Fréchet Inception Distance (FID) [13], which is applied to 2D rendered images of the objects.
4.2 Main Results
Quantitative Results. The quantitative results for the ShapeNet chairs and cars are presented in Table 1. Our observations are as follows: (1) In comparison to OP3D [12], which utilizes a single triplane representation to encode both shape and texture, GET3D [11] and our SYM3D, which separate shape and texture learning, exhibit improved generation in quality and diversity. (2) Models trained on various splits show that S1 outperforms S2, which in turn outperforms S3. Specifically, S1 encompasses a full 360-degree view of an object category, S2 achieves a 360-degree view through image flipping, while S3 is limited to a 240-degree view. The results further suggest that incorporating views from various angles, including those from other objects, enhances the learning of a 3D representation. (3) All models are trained without conditioning discriminator on camera poses. Our approach, which incorporates symmetry assumed in a canonical space, yields superior results in both COV and MMD metrics. This indicates that the shapes learned by SYM3D closely resemble those in a canonicalized form. (4) SYM3D outperforms GET3D in majority metrics, demonstrating the benefits of integrating symmetry regularization and view-wise spatial attention into 3D representations. Consequently, SYM3D proves to be an effective approach that enhances 3D representation.
Qualitative Comparisons. Figure 4 provides qualitative comparisons against competitors in terms of generated 2D images. In general, SYM3D achieves a more realistic appearance across different settings. In the more challenging settings S2, S3, and real-world ABO-chair, OP3D struggles to generate a complete scene, and GET3D tends to generate distortions. Figure 5 provides comparisons across various camera positions. GET3D and SYM3D are able to generate view-consistent images across different camera views but OP3D fails to do that. This demonstrates that decoupling shape and texture benefits the 3D-aware generation.
@@ 4.3 Properties of Learned Triplane
View-wise Triplanes. We use a similarity matrix to measure difference among feature maps in the geometry triplane. The geometry triplane is a $N\times N\times 3C$ feature tensor, which we flatten to a 2D tensor $3C\times N^{2}$.
@@
Robustness to Biased Views. Figure 1 shows the results from models trained on the incomplete view dataset: ShapNet Chair-S3. Overall, SYM3D shows its superiority in generating complete objects.
@@
Consistent Camera Orientations. The fixed camera distribution strategy ensures that all objects produced by both GET3D and SYM3D are uniformly positioned and oriented. By applying symmetry regularization, our method secures a consistent orientation for objects in the chair category with respect to the symmetry plane (Figure 8).
4.4 Further Analysis
The effect of different components. We validate the design of our framework by ablating four components using the ShapeNet Chair-S1 dataset. SVE introduces view-wise spatial enhancement in Section 3.2, $\mathcal{R}(G)$ adds symmetry in feature maps as in Eq. 5, $\mathcal{R}(A)$ adds symmetry in attention maps as in Eq. 6, $\mathcal{R}(T)$ supplements a point’s feature with its symmetric counterpart as in Eq. 7. Our findings indicate that each component independently improves the quality of generation, with the collective implementation of all elements resulting in the most effective model.
SVE $\mathcal{R}(G)$ $\mathcal{R}(A)$ Tex FID
@@ 4.4 Further Analysis
Table 2: ablation studies
With window self-attention. To show the effectiveness of the proposed attention module, we compare SYM3D against a variant that uses window self-attention [18] across three views, without employing attention symmetry. Considering memory limitations, we chose two window sizes: Wins=4 and Wins=8. Table 3 indicates that SYM3D, while adding negligible increases in parameters and computational cost, results in significant improvements.
@@
On a more complex category. For evaluating the generality of proposed method on more complex categories, we test on airplane dataset in ShapeNet [7]. We render 24 views for each object and select one view as training set. As shown in Table 4, we obtain consistent results as on ShapeNet-Chair and ShapeNet-Car in Table 1.
@@
Applied on the text-to-3D task. Current text-to-3D task usually utilize text-to-image generation models as guidance. The training process distills one image at a time, facing challenges in maintaining consistent views and precise geometry, leading to texture misalignments, asymmetry, incoherent appearances, or severe “Janus effect” issues where features like faces or eyes appear repeatedly and unnaturally [3] in generated object. Symmetry assumption encodes information from unseen views, providing a global constraint that defines a canonical frame, crucial for accurate model alignment and generation.
To demonstrate the generalizability of our method, we have applied our proposed symmetry regularization to this task. MTN [28] introduces a multi-scale triplane representation for the text-to-3D task. We apply symmetric regularization (see Eq 5) to these multi-scale triplanes. We develop two versions: W/O SYM and W SYM, whose results are in Figure 9. Our observations show that W/O SYM often generates salient artifacts, such as extra or malformed legs and ears when modeling a cat, and a dog with three forelegs. Conversely, W SYM produces cats and dogs without these defects, proving that symmetric priors effectively eliminate such artifacts in 3D model creation.
Figure 9: Comparisons on text-to-3D. W/O SYM produces artifacts in rendered images, while W SYM with symmetry regularization loss avoids this.
@@ 5 Conclusion
We present SYM3D for learning symmetric triplanes for improving 3D-awareness of GANs when trained on single images without camera pose annotation. While effective it has several limitations that future work should investigate. Global reflectional symmetry is not always satisfied and some objects satisfy other geometric transformations symmetry such as reflections, translations, rotations, or combinations in local parts (i.e. car tires, chair legs). In future work, we plan to extend our approach to create a large real dataset of common object categories and combat the canonicalization issue as in [2] and symmetry issue as in [20].